OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Sn: 1^2 + 4^2 + 7^2 + . . . + (3n - 2^)2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

Compute for n = 1, 2 and 3. each will be a separate case. show that the left side is equivalent to the right side for each case.

OpenStudy (anonymous):

S1 (LHS) = 1^2 = 1 S1 (RHS) = 1(6(1)^2 - 3(1) -1)/2 = (6-3-1)/2 = 2/2 = 1

OpenStudy (anonymous):

@pgpilot326 i honestly have no idea how to do this. can you break it down as much as you can please?